OpenStudy (anonymous):

cos300-sin(90-B) --------------- cos(-B)+sin330 how would this be solved using identities

OpenStudy (anonymous):

please anyone

ganeshie8 (ganeshie8):

use these : sin(90-B) = cosB cos(-B) = cosB sin(-B) = -sinB cos(x + 360) = cos(x) sin(x + 360) = sin(x)

OpenStudy (anonymous):

i have but i get stuck here \[[(0.5)-cosB]/[cosb-1/2]\]

ganeshie8 (ganeshie8):

wonderful! one more step you're done !!

OpenStudy (dan815):

ONE MOAR

OpenStudy (dan815):

ROW ROOW ROOWW YOUR BOAT

OpenStudy (dan815):

GENTLY DOWWNN THE STREAM

ganeshie8 (ganeshie8):

lol dan

OpenStudy (dan815):

moral support

ganeshie8 (ganeshie8):

\( [(0.5)-cosB]/[cosb-1/2] \) \(-1[cosB - 0.5]/[cosb-1/2] \)

OpenStudy (anonymous):

is that factorisation

ganeshie8 (ganeshie8):

no, i just pulled -1 out

ganeshie8 (ganeshie8):

(a-b) = -(b-a)

OpenStudy (anonymous):

ooh then you can cross cancel

ganeshie8 (ganeshie8):

exactly !

OpenStudy (anonymous):

ooh thank you so much

ganeshie8 (ganeshie8):

np :)